OpenStudy (anonymous):

use the distributive property to write the expression without parentheses 8(x+y)=

OpenStudy (anonymous):

i came up with 8x times 8y but im not sure or either 4x*2y

OpenStudy (anonymous):

@Jhannybean @zzr0ck3r

OpenStudy (zzr0ck3r):

8x + 8y

OpenStudy (jhannybean):

You just multiply the integer 8 to everything inside the parenthesis. Exactly what @zzr0ck3r showed.

OpenStudy (zzr0ck3r):

multiplication is distributive OVER addition.

OpenStudy (zzr0ck3r):

a(b+c) =ab+ac a(b+c+d) = ab+ac+ad a(b+c+d+e) = ab+ac+ad+ae . . .